OpenStudy (anonymous):

prove the identity: sin theta + (cos theta/tan theta)=1/(cos theta tan theta)

OpenStudy (raden):

sinθ + (cosθ/tanθ) = (sinθ * tanθ/tanθ) + (cosθ/tanθ) = (sinθ * tanθ + cosθ)/tanθ = (sinθ * sinθ/cosθ + cosθ)/tanθ = (sin^2 θ/cosθ + cosθ * cosθ/cosθ)/tanθ = (sin^2 θ/cosθ + cos^2 θ/cosθ)/tanθ = ( [sin^2 θ+cos^2 θ]/cosθ)/tanθ ) = (1/cosθ)/tanθ = 1/(cosθ*tanθ) LHS = RHS